OpenStudy (anonymous):

Jimmy's Chocolate shop sells two different fudge bars. The dark chocolate fudge bar sells for $6.00 per pound and the milk chocolate fudge bar sells for $4.00 per pound. You purchased a combination of these bars for all of your family and friends for the holidays and your bill came to $120. This scenario can be modeled by the equation 6d + 4m = 120. Which of the following is not a possible combination of dark chocolate bars and milk chocolate bars? 10 dark bars and 15 milk bars 12 dark bars and 12 milk bars 14 dark bars and 18 milk bars 6 dark bars and 21 milk bars

OpenStudy (tkhunny):

6(10) + 4(15) = 60 + 60 = 120 -- That one looks good. You do the next one.

OpenStudy (anonymous):

Thank you, you deserve a medal :-)